Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Selecting the color blind person in no more than 3 tries

Expert replies
by gmattesttaker2 » Tue May 27, 2014 6:32 pm
Hello,

For the following:

A randomly selected sample population consists of 60% women and 40% men. 90% of the women and 15% of the men are colorblind. For a certain experiment, scientists will select one person at a time until they have a colorblind subject. What is the approximate probability of selecting a colorblind person in no more than three tries?

A) 95%
B) 90%
C) 80%
D) 75%
E) 60%

OA: [spoiler]95%[/spoiler]


I was able to come up with the double set matrix but then couldn't solve for the probability part. This is what I have:

60/100 + (40/100).(60/99) + (40/100)(39/99)(60/98)

I wasn't sure though how to simplify this further. Thanks for your help.

Regards,
Sri
You do not have the required permissions to view the files attached to this post.
Join the discussion
Source: — Problem Solving |

by [email protected] » Tue May 27, 2014 9:38 pm
Hi Sri,

There are a couple of different ways to solve this problem. Here's how I solved it.

Based on the given info, we have….

60% Women
40% Men

Color-Blind:
90% of Women = (.9)(60%) = 54% of total group
15% of Men = (.15)(40%) = 6% of the total group

NOT Color-Blind = everyone else = 40% of the total group

Probability Rule:

Want + Don't Want = 1

When a probability question gives us multiple tries to accomplish 1 task, it's usually fastest to calculate the probability that we DO NOT accomplish the task, then subtract that from the number 1.

Instead of calculating the probability of picking a color-blind person in 3 tries, we'll calculate the probability that we DO NOT select a color-blind person in 3 tries….

Notice the word "approximate"….we can keep things simple….

(NOT)(NOT)(NOT) = (.4)(.4)(.4) = .064 This is the probability of NOT selecting a color-blind person in 3 tries.

The probability of selecting a color-blind person in 3 tries is….

1 - .064 = .936

This is really close to….

Final Answer: A

GMAT assassins aren't born, they're made,
Rich
Contact Rich at [email protected]
Image
Join the discussion

by buddyme » Tue Sep 22, 2015 1:32 am
[email protected] wrote:Hi Sri,

There are a couple of different ways to solve this problem. Here's how I solved it.

Based on the given info, we have….

60% Women
40% Men

Color-Blind:
90% of Women = (.9)(60%) = 54% of total group
15% of Men = (.15)(40%) = 6% of the total group

NOT Color-Blind = everyone else = 40% of the total group

Probability Rule:

Want + Don't Want = 1

When a probability question gives us multiple tries to accomplish 1 task, it's usually fastest to calculate the probability that we DO NOT accomplish the task, then subtract that from the number 1.

Instead of calculating the probability of picking a color-blind person in 3 tries, we'll calculate the probability that we DO NOT select a color-blind person in 3 tries….

Notice the word "approximate"….we can keep things simple….

(NOT)(NOT)(NOT) = (.4)(.4)(.4) = .064 This is the probability of NOT selecting a color-blind person in 3 tries.

The probability of selecting a color-blind person in 3 tries is….

1 - .064 = .936

This is really close to….

Final Answer: A

GMAT assassins aren't born, they're made,
Rich
Hi Rich, Is 0.936 the exact answer or is it an approximaation?
My doubt is that even if you have to calculate the probability that we DO NOT select a color-blind person in 3 tries, according to my undestanding, it hs to befollowing

40/100 * 39/99 * 38/98 and not 0.4 * 0.4 * 0.4 (since the selected candidate will not be placed back again)
If this is wrong, can you explain why is it wrong?

Actually if i take the total sample to be 200 instead of 100 then my answer will be different according to my calculations. So I am suspecting that the question itself is incorrect
Join the discussion

by [email protected] » Tue Sep 22, 2015 8:24 am
Hi buddyme,

Notice the wording of the question: "what is the APPROXIMATE probability..." - this means that whatever calculation you do will probably not be an exact match for any of the answer choices, so some 'rounding' will be required.

GMAT assassins aren't born, they're made,
Rich
Contact Rich at [email protected]
Image
Join the discussion

by Matt@VeritasPrep » Thu Sep 24, 2015 10:22 am
Let's call our population x.

.9*.6x + .4*.15x are colorblind, so .6x are colorblind.

The probability of selecting a colorblind person -- I assume they mean "at least one" colorblind person -- is just 1 - (probability of NO colorblind people), which is easier to compute.

Prob(no colorblind) = (first not colorblind) * (second not colorblind) * (third not colorblind)

They don't specify whether we're randomly selecting three different people (another big ambiguity), but let's assume that's NOT the case. We thus have

Prob(no colorblind) = .4x * .4x * .4x = .064x

So our probability of selecting a colorblind person is x - .064x, or .936x. No approximation required ... but the question is so ambiguously worded that we can't be perfectly sure 93.6% is the exact answer.
Join the discussion